National Income Accounts.
X 100. Transfers 100. PIT 220. SST 60.
IBT1 40. NPI 40. I 100. CRE 20. C 700.
NFFI 100. Depreciation 60. IBT2 40. G 300. M 200.
Determine NI
1
Expert's answer
2016-05-17T12:39:03-0400
The national income is defined as follows:
NI =GNP-depreciation-IBT1
Where
GNP=GDP+NFFI=C+I+G+(X-M)+NFFI
GNP = 700 + 100 + 300 + (100-200) +100 =1100
NI =1100 – 60 – 40 =1000